# p-MoD: Building Mixture-of-Depths MLLMs via Progressive Ratio Decay

### v1

@@ Abstract
Despite the remarkable performance of multimodal large language models (MLLMs) across diverse tasks, the substantial training and inference costs impede their advancement. The majority of computation stems from the overwhelming volume of vision tokens processed by the transformer decoder. In this paper, we propose to build efficient MLLMs by leveraging the Mixture-of-Depths (MoD) mechanism, where each transformer decoder layer selects essential vision tokens to process while skipping redundant ones. However, integrating MoD into MLLMs is non-trivial. To address the challenges of training and inference stability as well as limited training data, we adapt the MoD module with two novel designs: tanh-gated weight normalization (TanhNorm) and symmetric token reweighting (STRing). Moreover, we observe that vision tokens exhibit higher redundancy in deeper layer and thus design a progressive ratio decay (PRD) strategy, which gradually reduces the token retention ratio layer by layer, employing a shifted cosine schedule. This crucial design fully unleashes the potential of MoD, significantly boosting the efficiency and performance of our models. To validate the effectiveness of our approach, we conduct extensive experiments with two baseline models across 14 benchmarks. Our model, p-MoD, matches or even surpasses the performance of the baseline models, with only 55.6% TFLOPs and 53.8% KV cache storage during inference, and 77.7% GPU hours during training.
* Equal contribution. ✉ Corresponding author (lmwang@nju.edu.cn).
@@ 1 Introduction
Figure 1: Comparison with the LLaVA-NeXT [24] baseline. Our model demonstrates comparable performance with the baseline model on 14 benchmarks across various domains, with 46.2% fewer KV cache storage and 44.4% fewer TFLOPs during inference.
Recently, both academia and industry have witnessed the rapid development of multimdodal large language models (MLLMs) [39, 20, 44, 13, 41, 1], which have demonstrated exceptional performance across various visual-language understanding tasks. This success has positioned MLLMs as one of the most promising areas of AI research.
Reflecting on the roadmap of MLLM development, pioneering works [25, 47, 23, 7] focused on taking a single image as input and resizing it to a fixed low resolution that aligns with the training setup of the vision encoder. Subsequently, to meet the diverse demands of real-world applications, increasing efforts have broadened the forms of visual inputs that MLLMs can support, including multiple high-resolution images and videos[20, 39, 44, 5, 38].
Current state-of-the-art MLLMs handle high-resolution images either by dividing the original image into multiple slices [44, 24, 5] which are independently processed by the vision encoder, or by using a stronger vision encoder with improved positional encoding which can process any image at its native resolution [27, 39, 1]. Consequently, when processing multiple high-resolution images or videos, the number of vision tokens increases dramatically, proportional to the number of pixels and the number of images or video frames. The overwhelming volume of vision tokens processed by the transformer decoder results in explosion of computational costs, which severely hampers further development and broader application of MLLMs. Designing more efficient MLLM architectures with minimal performance degradation has thus become an urgent challenge for the community.
Previous efforts primarily focus on compressing vision tokens before the decoder with a lightweight connector [5, 44, 33, 40, 2]. These approaches forces the LLM to handle heavily compressed vision information, overlooking the fact that the LLM, with its enormous model capacity, has the potential to compress the vision tokens by itself within the transformer layer.
In this paper, we focus on optimizing computation efficiency of MLLMs within the transformer decoder layers. We draw inspiration from the Mixture-of-Depths (MoD[35] mechanism, originally developed for LLMs. Specifically, we propose to build efficient MLLMs with MoD mechanism by selecting the most important and informative vision tokens to be processed by each transformer layer, while skipping redundant ones to improve efficiency. However, integrating MoD mechanism into MLLMs is non-trivial and entails several significant challenges.
@@
Building upon the above innovations, we present our model, p-MoD, Mixture-of-Depth MLLMs equipped with our progressive ratio decay strategy and upgraded MoD layers (i.e. p-MoD layers). Extensive experiments validate the effectiveness of our proposed model. As shown in Figure 1, across 14 benchmarks spanning various domains, our model matches or even outperforms the strong LLaVA-NeXT [24] baseline, with only 55.6% TFLOPs and 53.8% KV cache storage during inference, and 77.7% GPU hours during training.
@@ 2 Related Work
Figure 2: Overview of p-MoD. In the left, we present the detailed architecture of our p-MoD layers. Given the input tokens, the weight predictor first assigns weights to each token. Then, a subset of tokens with highest weights are selected according to a pre-defined token retention ratio $R$ and processed by the transformer layer, while the rest of the tokens are skipped. The weights are normalized by TanhNorm module, and then both the selected and skipped tokens are symmetrically scaled by their corresponding weights in our STRing module. In the right, we demonstrate our crucial design, the progressive ratio decay(PRD) strategy, which gradually reduces the token retention ratio $R$ layer by layer, following a shifted cos schedule.
@@
Building Efficient MLLMs. The main computational burden in MLLMs arises from the large number of visual tokens that the transformer decoder must process. Previous works mainly focused on compressing visual tokens before they enter the decoder via convolution [5, 44], pooling [33, 40], or attention-based [2] (e.g. QFormer [21]) modules. However, reducing vision-related computation within the transformer decoder layers is less explored. LLaVolta [3] speeds up MLLM training by performing average pooling operation in intermediate transformer decoder layers. FastV [4] and some subsequent works [12, 45, 14] improve MLLMs’ efficiency by dropping vision tokens in intermediate transformer layers based on the attention scores they received from text tokens. However, these methods require the acquisition of the attention score matrix, making them incompatible with widely used efficient attention implementations like FlashAttention [9, 8] which significantly improves speed and memory efficiency during training and inference. Moreover, these methods neglect the fact that tokens discarded in early layers due to low attention scores might be crucial in deeper layers, potentially leading to a degradation in performance. In contrast, we utilize MoD module to perform layerwise vision token selection without relying on text, which not only enjoys the efficiency of FlashAttention, but also incurs negligible performance loss. We compare our method with these attention-score-based pruning methods in Section A in the supplementary material.
@@ 3.1 Revisiting Mixture-of-Depths
The MoD layer consists of a weight predictor and a vanilla transformer layer. It predicts an importance score for each input token using the weight predictor and then selects the top $R\%$ most important tokens to be processed by the transformer layer. Formally, given an MoD layer $M$ with a token retention ratio $R$, a transformer layer $T$, and a linear weight predictor, the input tokens $X\in\mathbb{R}^{n\times d}$ is first passed through the linear predictor to generate a set of weights:
$w=\text{Linear}(X)\in\mathbb{R}^{n}$ (1)
@@
Then, we compute the $R$-th percentile of the router weights, denoted as $P_{R}(w)$. Tokens with weights larger than $P_{R}(w)$ will be selected to be processed by the transformer layer $T$ while other tokens are skipped in this layer, which guarantees only $n\times R\%$ tokens are processed. The calculation process in transformer layer $T$ with MoD can be formulated as:
$X_{i}^{\prime}=\begin{cases}w_{i}T(X_{i})+X_{i},&\text{if}\quad w_{i}>P_{R}(w)\\ X_{i},&\text{if}\quad w_{i}\leq P_{R}(w)\end{cases}$ (2)
@@
In the original MoD module designed for LLMs [35], $X$ represents text tokens. In our multimodal scenario, $X$ represents vision tokens. We only apply MoD on vision tokens as they occupy the main computation load and exhibit high redundancy in deep layers.
@@ 3.2.1 Tanh-gated Weight Normalization
$X_{i}^{\prime}=\begin{cases}\alpha\tanh(w_{i})T(X_{i})+X_{i},&\text{if}\quad w_{i}>P_{R}(w)\\ X_{i},&\text{if}\quad w_{i}\leq P_{R}(w)\end{cases}$ (3)
@@ 3.2.2 Symmetric Token Reweighting
$X_{i}^{\prime}=\begin{cases}\alpha\tanh(w_{i})T(X_{i})+X_{i},&\text{if}\quad w_{i}>P_{R}(w)\\ \alpha\tanh(w_{i})X_{i}+X_{i},&\text{if}\quad w_{i}\leq P_{R}(w)\end{cases}$ (4)
@@ 3.3 Progressive Ratio Decay
(a) Illustration of our exploratory experiments on our MoD-based MLLM trained with a token retention ratio of 70% across all layers. Each sub-figure shows the inference token retention ratio across all layers of one experiment. Experiment $a$ is the baseline setting (70% ratio). In experiments $b$, $c$ and $d$, we apply a lower token retention ratio (15%) to shallow, middle and deep layers.
(b) Average performance of our four experiments on 14 benchmarks listed in Section 4.1. The performance drop becomes less significant as the low token retention ratio is applied to deeper layers, which indicates that vision tokens exhibit higher redundancy in deeper layers.
Figure 3: Exploratory experiments on vision token redundancy.
@@
To validate our hypothesis, we conduct a series of exploratory experiments on our MoD-based MLLM trained with a token retention ratio of 70% across all layers. As shown in Figure 3, we first divide the layers into several groups. In each experiment, we apply a lower token retention ratio (15%) only to one group of layers while keeping the other layers unchanged, and evaluate the model under this customized inference setting on 14 benchmarks. The results in Figure 3 strongly support our hypothesis: the deeper the low token retention ratio is applied, the better the performance. This suggests that vision tokens exhibit higher redundancy in deeper layers, which is consistent with observations in previous works [4, 45].
Accordingly, we design a progressive ratio decay (PRD) strategy. As shown on the right side of Figure 2, PRD reduces the token retention ratio layer by layer, following a shifted cosine schedule. Suppose the model has $L$ layers, the token retention ratio for the $l$-th layer is formulated as:
$R_{l}=\cfrac{1}{2}\cos{\cfrac{\pi l}{L}}+\beta,\quad l=1,2,...,L$ (5)
@@
Model Params
Inference
@@
Inference KV
cache storage
SEED RWQA MME MMB POPE GQA AI2D AVG
LLaVA-v1.5 7B 8.38 100% 66.2 55.6 1506.8 64.1 85.9 61.9 55.2 66.3
p-MoD-LLaVA-v1.5 7B 4.92${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-41.3\%)}$ 53.8% 66.5 55.7 1482.8 65.4 85.5 62.2 56.2 66.5
LLaVA-NeXT 7B 39.46 100% 68.9 57.6 1519.3 67.5 87.2 63.5 64.0 69.2
p-MoD-LLaVA-NeXT 7B 21.94${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(-44.4\%)}$ 53.8% 69.0 57.6 1495.5 67.3 86.8 63.3 65.1 69.1
Table 1: Comparison with baseline models on 7 benchmarks. p-MoD achieves comparable or even better performance compared to baseline models, with significantly improved efficiency. Params stands for the number of model parameters. We also report the Inference TFLOPs and KV cache storage savings to compare the efficiency of p-MoD and the baseline models. Benchmark names are abbreviated. SEED: SEED-Bench (Image); RWQA: RealWorldQA; MME: MME Perception; MMB: MMBench. AVG stands for the average performance in percentage across all benchmarks. MME Perception score is divided by 2000 to compute the average. We set the gating factor $\alpha$ in TanhNorm to 0.2, and the shift factor $\beta$ in PRD to $0.5$.
Model Params DocVQA ChartQA TextVQA InfoVQA OK-VQA MMMU SQA AVG
LLaVA-v1.5 7B 28.1 18.2 46.0 25.8 53.4 36.6 69.7 39.7
p-MoD-LLaVA-v1.5 7B 27.6 16.8 44.8 26.8 56.0 36.3 69.3 39.7
LLaVA-NeXT 7B 70.1 61.6 62.7 34.7 54.0 35.3 69.7 55.5
p-MoD-LLaVA-NeXT 7B 70.0 61.8 60.5 34.1 55.1 36.0 71.0 55.5
Table 2: Comparison with baseline models on other 7 benchmarks. Our p-MoD models matches the performance of corresponding baseline models. SQA stands for ScienceQA-IMG. We set the gating factor $\alpha$ in TanhNorm to 0.2, and the shift factor $\beta$ in PRD to $0.5$.

### v2

@@ Abstract
Despite the remarkable performance of multimodal large language models (MLLMs) across diverse tasks, the substantial training and inference costs impede their advancement. In this paper, we propose p-MoD, an efficient MLLM architecture that significantly reduces training and inference costs while maintaining model performance. The majority of computation in MLLMs stems from the overwhelming volume of vision tokens processed by the transformer-based LLM. Accordingly, we leverage the Mixture-of-Depths (MoD) mechanism, where each LLM layer selects essential vision tokens to process while skipping redundant ones. However, integrating MoD into MLLMs is non-trivial. To address the challenges of training and inference stability as well as limited training data, we adapt the MoD module with two novel designs: tanh-gated weight normalization (TanhNorm) and symmetric token reweighting (STRing). Moreover, we observe that vision tokens exhibit higher redundancy in deeper layers and thus design a progressive ratio decay (PRD) strategy, which gradually reduces the token retention ratio layer by layer, employing a shifted cosine schedule. This crucial design fully unleashes the potential of MoD, significantly boosting the efficiency and performance of our models. Extensive experiments on two baseline models across 15 benchmarks show that our model matches or even surpasses the performance of corresponding baselines, while requiring only 55.6% TFLOPs and 53.7% KV cache storage during inference, and 77.7% GPU hours during training.
00footnotetext: * Equal contribution. 🖂 Corresponding author.
@@ 1 Introduction
Figure 1: Comparison with the LLaVA-NeXT [28] baseline. Our model demonstrates comparable performance with the baseline model on 15 benchmarks across various domains, with 46.3% fewer KV cache storage and 44.4% fewer TFLOPs during inference.
Recently, both academia and industry have witnessed the rapid development of multimdodal large language models (MLLMs) [47, 24, 55, 14, 51, 1, 46], which have demonstrated exceptional performance across various vision-language understanding tasks. Pioneering works in this field [29, 59, 27, 8] focused on processing single low-resolution image inputs. Subsequently, to meet the diverse demands of real-world applications, increasing efforts have broadened the forms of visual inputs that MLLMs can support, including multiple high-resolution images and videos [24, 47, 55, 6, 45].
Current state-of-the-art MLLMs handle high-resolution images either by dividing the original image into multiple slices [55, 28, 6] which are independently processed by the vision encoder, or by using a stronger vision encoder with improved positional encoding which can process any image at its native resolution [31, 47, 1]. Consequently, when processing multiple high-resolution images or videos, the number of vision tokens increases dramatically, proportional to the number of pixels and the number of images or video frames. The overwhelming volume of vision tokens processed by the transformer-based LLM results in explosion of computational costs, which severely hampers further development and broader application of MLLMs. Designing more efficient MLLM architectures with minimal performance degradation has thus become an urgent challenge for the community.
Previous efforts primarily focus on compressing vision tokens before the LLM, either in the Vision Encoder or the multimodal projector [6, 55, 38, 50, 2, 12, 42, 49, 56, 39]. These approaches forces the LLM to handle heavily compressed vision information, overlooking the fact that the LLM, with its enormous model capacity, has the potential to compress the vision tokens by itself within the transformer layer.
In this paper, we focus on optimizing computation efficiency of MLLMs within the transformer layers of the LLM. We propose to build efficient MLLMs with Mixture-of-Depths (MoD[41] mechanism, which selects the most important and informative vision tokens to be processed by each transformer layer, while skipping redundant ones to improve efficiency. However, integrating MoD mechanism into MLLMs is non-trivial and entails several significant challenges.
@@
Building upon the above innovations, we present our model, p-MoD, Mixture-of-Depth MLLMs equipped with our progressive ratio decay strategy and upgraded MoD layers (i.e. p-MoD layers). Extensive experiments validate the effectiveness of our proposed model. As shown in Figure 1, across 15 benchmarks spanning various domains, our model matches or even outperforms the strong LLaVA-NeXT [28] baseline, with only 55.6% TFLOPs and 53.7% KV cache storage during inference, and 77.7% GPU hours during training.
@@ 2 Related Work
Figure 2: Overview of p-MoD.  Middle: Our efficient MLLM model which consists of p-MoD layers. Each layer independently selects important vision tokens to process while skipping redundant ones. The proportion of selected tokens for layer $i$ is specified by the token retention ratio $R_{i}$. For simplicity, the vision encoder and connector are omitted.  Left: Detailed architecture of p-MoD layers. Given the input tokens, the weight predictor first assigns weights to each token. The top $R_{i}$% of tokens with highest weights are selected and processed by the transformer layer, while the rest of the tokens are skipped. The weights are normalized by TanhNorm module, and then both the selected and skipped tokens are symmetrically scaled by their corresponding weights in our STRing module.  Right: Our crucial design is the progressive ratio decay(PRD) strategy, which gradually reduces the token retention ratio $R_{i}$ layer by layer, following a shifted cos schedule.
@@
Building Efficient MLLMs. The main computational burden in MLLMs arises from the large number of vision tokens that the LLM must process. Previous works mainly focused on compressing vision tokens before they enter the LLM via convolution [6, 55], pooling [38, 50], query-based modules [2, 16] or token merging [3, 25, 49, 42, 15]. However, reducing vision-related computation within the LLM layers is less explored. LLaVolta [4] speeds up MLLM training by performing naive average pooling operation in intermediate LLM layers. FastV [5] and some subsequent works [13, 57, 17, 48, 15] improve MLLMs’ efficiency by compressing vision tokens in intermediate transformer layers based on attention scores. However, these methods neglect the fact that tokens compressed in early layers might be crucial in deeper layers, potentially leading to performance degradation. In contrast, we utilize MoD for learnable layerwise vision token selection, improving model efficiency while avoiding potential performance loss caused by token dropping.
@@ 3.1 Revisiting Mixture-of-Depths
The MoD layer consists of a weight predictor and a vanilla transformer layer. It assigns weights for each input token using the weight predictor and then selects the top $R\%$ tokens with the highest weights to be processed by the transformer layer. Formally, given an MoD layer $M$ with a token retention ratio $R$, a transformer layer $T$, and a linear weight predictor, the input tokens $X\in\mathbb{R}^{n\times d}$ is first passed through the linear predictor to generate a set of weights:
$w=\text{Linear}(X)\in\mathbb{R}^{n},$ (1)
@@
Then, we compute the $R$-th percentile of the router weights, denoted as $P_{R}(w)$. Tokens with weights larger than $P_{R}(w)$ will be selected to be processed by the transformer layer $T$ while other tokens are skipped in this layer, which guarantees only $n\times R\%$ tokens are processed. The calculation process of an MoD layer can be formulated as:
$X_{i}^{\prime}=\begin{cases}w_{i}T(X_{i})+T(X_{i}),&\text{if}\quad w_{i}>P_{R}(w)\\ X_{i},&\text{if}\quad w_{i}\leq P_{R}(w)\end{cases}.$ (2)
@@
In the original MoD module designed for LLMs [41], $X$ represents text tokens. In our multimodal scenario, $X$ represents vision tokens. We only apply MoD on vision tokens as they occupy the main computation load and exhibit high redundancy in LLM layers.
@@ 3.2 Adapting Mixture-of-Depths Module
(a) Illustration of our exploratory experiment setup. Our baseline MoD model is trained with a token retention ratio of 70% for all layers. We divide the layers into three groups: shallow, middle, and deep. In each set of experiments, we decrease the token retention ratio (R%) in a specific group of layers.
(b) Results of our exploratory experiments. Performance drops more slowly when the token retention ratio of deeper layers is decreased, indicating that vision tokens are more redundant in deeper layers. This inspires us to progressively reduce the token retention ratio layer by layer.
Figure 3: Exploratory experiments on vision token redundancy.
@@ 3.2.1 Tanh-gated Weight Normalization
$X_{i}^{\prime}=\begin{cases}\alpha\tanh(w_{i})T(X_{i})&\\ \quad\quad\quad\quad+T(X_{i}),&\text{if}\quad w_{i}>P_{R}(w)\\ X_{i},&\text{if}\quad w_{i}\leq P_{R}(w)\end{cases}\quad.$ (3)
@@
Model
Inference
TFLOPs $\downarrow$
Inference
KV cache $\downarrow$
Doc
VQA
Chart
QA
Text
VQA
Info
VQA
RW
QA
G
QA
OK
VQA
PO
PE
AI
2D
SE
ED
AVG
LLaVA-v1.5 8.38 100% 28.1 18.2 46.0 25.8 55.6 61.9 53.4 85.9 55.2 66.2 49.6
[…]
LLaVA-NeXT 39.46 100% 70.1 61.6 62.7 34.7 57.6 63.5 54.0 87.2 64.0 68.9 62.4
[…]
Table 1: Comparison with baseline models on 10 benchmarks. Our models (marked in gray) achieves comparable or even better performance compared to baseline models, with only 55.6% TFLOPs and 53.7% KV cache storage during inference. All models are of 7B parameter scale.
Model
S
QA
$M$
B
$M$
MU
VQA
v2
$M$
E
AVG
LLaVA-v1.5 69.7 64.1 36.6 76.6 1506.8 64.5
+ p-MoD 69.3 65.4 36.3 76.9 1482.8 64.4
LLaVA-NeXT 69.7 67.5 35.3 79.3 1519.3 65.6
+ p-MoD 71.0 67.3 36.0 78.8 1495.5 65.6
Table 2: Comparison with baseline models on more benchmarks. Our p-MoD models (marked in gray) matches the performance of corresponding baseline models. SQA stands for ScienceQA-IMG. All models are 7B in size.
Model
Doc
VQA
Chart
QA
Text
VQA
G
QA
SE
ED
$M$
B
AVG
vanilla MoD 61.1 54.0 55.8 61.8 66.9 63.1 60.4
+ TanhNorm 61.8 54.5 56.4 62.5 67.1 65.9 61.4
   + STRing 65.7 58.4 59.3 63.0 67.1 66.8 63.4
    + PRD 70.0 61.8 60.5 63.3 69.0 67.3 65.3
Table 3: Ablation study on our propsed innovations. Under fair comparison, all of our proposed modules (TanhNorm, STRing, and PRD) significantly improve model performance without any computational overhead, making indispensable contributions to the strong performance of our final model (marked in gray).
@@ 3.2.2 Symmetric Token Reweighting
$X_{i}^{\prime}=\begin{cases}\alpha\tanh(w_{i})T(X_{i})&\\ \quad\quad\quad\quad+T(X_{i}),&\text{if}\quad w_{i}>P_{R}(w)\\ \alpha\tanh(w_{i})X_{i}+X_{i},&\text{if}\quad w_{i}\leq P_{R}(w)\end{cases}\quad.$ (4)
@@ 3.3 Progressive Ratio Decay
To validate our hypothesis, we conduct a series of exploratory experiments on our MoD-based MLLM trained with a token retention ratio of 70% across all layers. As shown in Figure 3(a), we first divide the layers into several groups: shallow, middle, and deep layers. In each set of experiment, we gradually decrease the token retention ratio (R%) in a specific group of layers while keeping the other layers unchanged, and evaluate the model under this customized inference setting. The results in Figure 3(b) strongly support our hypothesis: performance drops more slowly when the token retention ratio of deeper layers is decreased. This suggests that vision tokens exhibit higher redundancy in deeper layers, which is consistent with observations in previous works [5, 57].
Accordingly, we design a progressive ratio decay (PRD) strategy. As shown on the right side of Figure 2, PRD gradually reduces the token retention ratio layer by layer, following a shifted cosine schedule. Suppose the model has $L$ layers, the token retention ratio for the $l$-th layer is formulated as:
$R_{l}=\cfrac{1}{2}\cos{\cfrac{\pi l}{L}}+\beta,\quad l=1,2,...,L.$ (5)
@@
Ratio
Scheduler
Decay Progressive
Doc
VQA
Chart
QA
Text
VQA
G
QA
SE
ED
$M$
B
AVG
Constant $usym{2717} \usym$$\usym{2717}$ 58.0 51.6 56.1 62.1 64.5 62.8 59.2
Interleaved $usym{2717} \usym$$\usym{2717}$ 60.3 54.8 57.0 62.3 65.3 62.1 60.3
Stepped $usym{2714} \usym$$\usym{2717}$ 67.2 61.2 59.4 63.5 68.4 66.1 64.3
Linear 69.3 60.4 60.5 63.8 68.0 66.7 64.8
PRD Cosine $usym{2714} \usym$$\usym{2714}$ 70.0 61.8 60.5 63.3 69.0 67.3 65.3
Model
Token Compression
Ratio $\downarrow$
Infer.
@@
Doc
VQA
Chart
QA
Text
VQA
Info
VQA
RW
QA
GQA SQA
$M$
MU
PO
PE
AI
2D
VQA
v2
SE
ED
AVG
LLaVA-NeXT 100% 39.46 70.1 61.6 62.7 34.7 57.6 63.5 69.7 35.3 87.2 64.0 79.3 68.9 62.9
 + MQT [16] 50.2% 19.86 49.9 44.0 53.5 29.8 53.5 62.3 69.1 36.1 85.8 64.0 77.2 65.8 57.6
 + LLaVolta [4] 53.1% 21.30 66.6 58.8 60.1 33.2 56.9 63.7 70.2 36.7 86.3 65.3 78.6 68.5 62.0
 + FastV [5] 53.1% 22.73 65.9 58.2 62.2 33.0 55.6 62.4 69.9 35.6 84.5 63.6 78.7 68.2 61.5
+ p-MoD 53.7% 21.94 70.0 61.8 60.5 34.1 57.6 63.3 71.0 36.0 86.8 65.1 78.8 69.0 62.8
Table 5: Comparison with other vision token compression methods. p-MoD significantly outperforms other compression methods, achieving the best performance on most benchmarks and the best average performance. All models are of 7B parameter scale. Results on 13B models are demonstrated in Table 8.
4 Experiment
4.1 Setups
Models. To evaluate the effectiveness of p-MoD, we select two representative open-source MLLMs: LLaVA-1.5 [27] and LLaVA-NeXT [28], as the baseline models in our experiments. LLaVA-1.5 resizes input images to the fixed resolution which aligns with the training setup of its CLIP [40] vision encoder, encoding an image into 576 tokens. LLaVA-NeXT divides high-resolution images into multiple slices which are independently processed by the vision encoder. This strategy enhances the visual perception capabilities, but also leads to a significantly larger number of vision tokens (up to 2880) and much higher computation costs.
Training Recipe. Our p-MoD models follow the two-stage training recipe of the baselines. In the pre-training stage, the MLP connector is trained on image caption data. In the fine-tuning stage, p-MoD modules are insert into the LLM and updated together with the connector and the LLM.
Benchmarks. We conduct comprehensive experiments across 15 benchmarks: VQAv2 [11], GQA [18], OK-VQA [34], AI2D [21] and ScienceQA-IMG [32] are traditional visual question answering benchmarks; DocVQA [36], TextVQA [44], ChartQA [35] and InfographicVQA [37] focus on fine-grained visual question answering; POPE [26] evaluates hallucination in MLLMs; SEED-Bench (Image) [23], RealWorldQA, MME [10], MMBench [30] and MMMU  [53] are comprehensive benchmarks tailored for MLLMs. To ensure our results can be conveniently reproduced, we evaluate our model on all these benchmarks with the lmms-eval [54] evaluation framework.
Evaluation Metrics. In addition to performance on benchmarks, we present various metrics that reflect training and inference efficiency. We report TFLOPs which measures the inference computation complexity, and the KV cache storage which constitutes the main memory bottleneck during inference. Furthermore, we measure GPU hours consumed during training, along with inference latency.
4.2 Main Results
We integrate p-MoD with LLaVA-1.5 and LLaVA-NeXT baseline models and conduct comprehensive evaluation across 15 benchmarks, which measures comprehensive abilities of MLLMs across various aspects. The results are shown in Table 1 and Table 3. Both p-MoD-LLaVA-v1.5 and p-MoD-LLaVA-NeXT achieve comparable or even better performance across all benchmarks compared to their baseline models, with significant savings in inference TFLOPs and inference KV cache savings.
4.3 Ablation Study
Effectiveness of propsed innovations. We conduct ablation in Table 3 to demonstrate the effectiveness of our proposed innovations. All experiments in the table are conducted under the same computational cost. TanhNorm, STRing and PRD all significantly enhance model performance, making indispensable contributions to the strong performance of our final p-MoD model.
To further validate this conclusion and the effectiveness of PRD, we experiment with four other schedule functions which control the token retention ratio for each LLM layer. As illustrated in Figure 4, the functions include: (1) a constant function that uses the same ratio for all layers, (2) an interleaved function which interleaves a vanilla transformer layer and an MoD layer with low ratio, which is adopted in the original MoD paper [41], (3) a stepped decay function, (4) a linear decay function. To ensure a fair comparison, we fix the average retention ratios across all layers at approximately 54%. Based on the results demonstrated in Table 4, we can draw the following conclusions:
  • Our progressively decayed PRD functions (i.e. cosine and linear) notably outperforms the discontinuous stepped decay function, and the cosine function achieves the best results. This proves the effectiveness of our shifted cosine schedule.
Tanh-gated Weight Normalization. To validate the effectiveness of TanhNorm, we compare different weight normalization methods in Table 6. First, we verify that using no weight normalization module will result in overflow during training, as shown in the first row (details are explained in appendix). Then, we validate the effectiveness of the two important properties of TanhNorm stated in Section 3.2.1: (1) the normalized weight distribution is zero-centered, (2) the range(variance) of the weight distribution can be easily controlled by adjusting the gating factor $\alpha$.
Norm
Type
Doc
VQA
Chart
QA
Text
VQA
G
QA
SE
ED
$M$
B
AVG
- OVERFLOW -
Softmax 63.9 57.6 57.9 62.7 68.2 67.0 62.9
Shifted softmax 62.5 57.0 57.3 62.9 68.3 66.4 62.4
[…]
Table 6: Ablation on tanh-gated weight normalization. Our default model is marked in gray. The rows with no performance reported indicates that the corresponding experiment faces overflow issue during training or inference.
4.4 Comparison with Related Works
In Table 5, we compare p-MoD with several strong token compression methods. MQT [16] employs a query transformer to compress vision tokens. LLaVolta [4] performs average pooling in intermediate LLM layers, achieving remarkable results despite its simplicity. FastV [5] drops vision tokens in intermediate LLM layers based on the attention score from text tokens. p-MoD significantly outperforms other compression methods, achieving the best performance on most benchmarks and the best average performance.